# Cross-Domain Federated Semantic Communication with Global Representation Alignment and Domain-Aware Aggregation

### v1

@@ IV-D Swin Transformer-based DeepJSCC
In this section, we provide the details of the DL architecture of the semantic communication model. Specifically, we implement the Swin Transformer model[liu2021swin], which has high capability in semantic extraction and also exhibits high computing efficiency, for the source encoder/decoder. Instead of executing a global attention mechanism over all the patches within the image, like the conventional Vision Transformer, the Swin model partitions the image into non-overlapping local windows and applies the attention mechanism to each window. Benefiting from this local self-attention, we can significantly reduce the computing operation by processing on a smaller region. To capture the long-range dependency correlation across windows, Swin Transformer proposes a shifted window operation to provide connections among windows. With an image having $h\times w$ patches, and each window contains $M\times M$ patches, the comparison of computation complexity between the global multi-head self-attention (MSA) and the window-based is presented according to [liu2021swin]
@@
The MSA scales quadratically with the number of patches $hw$, and W-MSA only scales linearly with the image size, while we can set the size of $M$. Specifically, our encoder model consists of two Swin Transformer stages, where the first stage comprises two Transformer blocks, and the subsequent stage features four blocks. On the other hand, the source decoder has a symmetric architecture toward the source encoder, where the first stage has four Transformer blocks and the second has two. The symmetric architecture between the source encoder and decoder enhances the capability for signal decoding, thereby improving end-to-end performance. On the other hand, we adopt a simple yet effective DL architecture [DJSCCSwin] for the channel coders to compress and decompress the semantic features. Specifically, it consists of a set of fully connected layers combined with an SNR fusion mechanism, which enables the model to adapt its encoding and decoding process according to the current channel condition, thereby improving robustness and reconstruction fidelity under varying noise levels.
@@ V Simulation Results and Analysis
Within this section, we conduct a comprehensive series of simulations to effectively demonstrate the outstanding performance of our frameworks when the client’s data originates from various data sources. First, we provide the simulation setting for our training framework and also the communication environment for the semantic communication system. Secondly, we present benchmark schemes for comparison analysis, and finally, we evaluate system performance under our proposed approaches in comparison to others. As our framework addresses domain shift in federated learning, we refer to it as FedDoM.
@@ V-A1 Datasets, Distributions, and DL-based architecture
We adopt the PACS dataset for our training and evaluation process due to its diverse image domains, including photo, sketch, art painting, and cartoon[li2017deeper]. Similar to other studies, we sample $30\%$ of the full-size datasets for each domain, and we divide them into $80\%$ for training process and $20\%$ for testing. Then, we partition these data samples into a total of ten clients using the Dirichlet distribution with $\alpha=1$. Specifically, for each domain, the data is divided into two or three clients, and each client is limited to one type of domain data, which exacerbates the heterogeneity in domains. We provide the number of data samples of clients and their domains as in Table I.
@@
We apply the power normalization operation to the output of the DL-based JSCC encoder before transmitting it, as described in [DJSCCnofeedback]. To simulate the dynamic of channel noise, we randomly sample a value from the set of $\mathcal{S}$ $\in\{1,3,5,7,9\}$, while the transmit power is fixed at $1$ for all the experiments. The learning rate $\eta$ of DL models is $1e^{-3}$, the training batch size is $16$, and the total global training rounds are $500$.
@@ V-A2 Benchmarks
  • FedAvg: It is a traditional federated learning framework, where clients are trained independently of each other, and the BS aggregates the global model using the proportion of the number of data samples from each client[mcmahan2017communication].
  • FedProx: It introduces a regularization term to limit the deviation of each local model from the global model in the local training process, thereby helping to mitigate data heterogeneity among clients in the network [li2020federated].
  • MOON: It proposes a contrastive loss that encourages each client’s model to stay closer to the global model and farther from outdated local models, helping align representations across clients[li2021model].
@@ V-B1 Performance Convergence
AWGN
Channel
Condition SNR
FedAvg FedProx MOON FedDoM (Ours)

### v2

@@ IV-D Swin Transformer-based DeepJSCC
In this section, we provide the details of the DL architecture of the semantic communication model. Specifically, we implement the Swin Transformer model[18], which has high capability in semantic extraction and also exhibits high computing efficiency, for the source encoder/decoder. Instead of executing a global attention mechanism over all the patches within the image, like the conventional Vision Transformer, the Swin model partitions the image into non-overlapping local windows and applies the attention mechanism to each window. Benefiting from this local self-attention, we can significantly reduce the computing operation by processing on a smaller region. To capture the long-range dependency correlation across windows, Swin Transformer proposes a shifted window operation to provide connections among windows. With an image having $h\times w$ patches, and each window contains $M\times M$ patches, the comparison of computation complexity between the global multi-head self-attention (MSA) and the window-based is presented according to [18]
@@
The MSA scales quadratically with the number of patches $hw$, and W-MSA only scales linearly with the image size, while we can set the size of $M$. Specifically, our encoder model consists of two Swin Transformer stages, where the first stage comprises two Transformer blocks, and the subsequent stage features four blocks. On the other hand, the source decoder has a symmetric architecture toward the source encoder, where the first stage has four Transformer blocks and the second has two. The symmetric architecture between the source encoder and decoder enhances the capability for signal decoding, thereby improving end-to-end performance. On the other hand, we adopt a simple yet effective DL architecture [43] for the channel coders to compress and decompress the semantic features. Specifically, it consists of a set of fully connected layers combined with an SNR fusion mechanism, which enables the model to adapt its encoding and decoding process according to the current channel condition, thereby improving robustness and reconstruction fidelity under varying noise levels.
@@ V Simulation Results and Analysis
Within this section, we conduct a comprehensive series of simulations to effectively demonstrate the outstanding performance of our frameworks when the client’s data originates from various data sources. First, we provide the simulation setting for our training framework and also the communication environment for the semantic communication system. Secondly, we present benchmark schemes for comparison analysis, and finally, we evaluate system performance under our proposed approaches in comparison to others. We refer to our framework as FedDoM.
@@ V-A1 Datasets, Distributions, and DL-based architecture
We adopt the PACS dataset for our training and evaluation process due to its diverse image domains, including photo, sketch, art painting, and cartoon[13]. Similar to other studies, we sample $30\%$ of the full-size datasets for each domain, and we divide them into $80\%$ for training process and $20\%$ for testing. Then, we partition these data samples into a total of ten clients using the Dirichlet distribution with $\alpha=1$. Specifically, for each domain, the data is divided into two or three clients, and each client is limited to one type of domain data, which exacerbates the heterogeneity in domains. We provide the number of data samples of clients and their domains as in Table I.
@@
We apply the power normalization operation to the output of the DL-based JSCC encoder before transmitting it, as described in [1]. To simulate the dynamic of channel noise, we randomly sample a value from the set of $\mathcal{S}$ $\in\{1,3,5,7,9\}$, while the transmit power is fixed at $1$ for all the experiments. The learning rate $\eta$ of DL models is $1e^{-3}$, the training batch size is $16$, and the total global training rounds are $500$.
@@ V-A2 Benchmarks
  • FedAvg: It is a traditional federated learning framework, where clients are trained independently of each other, and the BS aggregates the global model using the proportion of the number of data samples from each client[19].
  • FedProx: It introduces a regularization term to limit the deviation of each local model from the global model in the local training process, thereby helping to mitigate data heterogeneity among clients in the network [16].
  • MOON: It proposes a contrastive loss that encourages each client’s model to stay closer to the global model and farther from outdated local models, helping align representations across clients[14].
@@ V-B1 Performance Convergence
AWGN Channel Condition SNR FedAvg FedProx MOON FedDoM (Ours)